(Philly Air) PhillyAir Inc. offers low cost air travel between Philadelphia and Atlantic City. Philly Air's invested capital is $5,000,000, corresponding to the investment in the two planes the company owns. Each of the two planes can carry 50 passengers. Each plane does 12 daily trips from Philadelphia to Atlantic City and 12 from Atlantic City to Phila- delphia. The price is $100 for each one-way ticket. The current load factor is 70 percent (ie., 35 seats are sold on the average flight). The annual cost of operating the service and Q6.3
running the business is $60,000,000 (including all costs, such as labor, fuel, marketing, gate fees, landing fees, maintenance, etc). The company operates 365 days a year a. Draw an ROIC (return on invested capital) tree for the company that incorporates all of the above information. b. What is the current ROIC? c. What is the minimum load factor at which the company breaks even? d. What load factor would the company have to achieve so that it obtained a 10 percentage- point increase in the ROIC (e.g. an ROIC increasing from 5 percent to 15 percent)?
Answer #1

a) ROIC diagram

Margin $61,320,000 Margin Revenue - Cost - $1,320,000 ROIC - Margin/ Cost Investment $60,000,000 Investment $5,000,000

Total trips in a day for 2 planes = 2*12 *2= 48 (2 planes and 12 trips each way)

Total seats on average sold = 35

Number of days operated = 365

Total revenue= Price/ticket * Number of days operated * Total seats on average sold * Total trips in a day = 100 * 365*35*48 = 61320000

Cost of running business = 60,000,000

Net Profit = 61320000-60,000,000 = 1320000

ROIC (Current) = NEt profit / Investment = 1320000/ 5000000 = 26.40%

c) For breakeven , cost = revenue.

Hence,Price/ticket * Number of days operated * Load factor (X) *Total seats * Total trips in a day = 60,000,000

= 100*365*48* X*50 = 60,000,000

87600000 X =  60,000,000

X = 68.49 %

d) NEW ROIC =  26.4 + 10 = 36.4% ROIC = Net profit / Investment

Hence, (100*365*48* X*50 - 60,000,000) / 5000000 = 36.4%

87600000 X - 60,000,000 = 36.4%*5000000

X = 70.57%